GreenHab Report – March 27th

Crew 207 GreenHab Report – 27-Mar-2019

GreenHab Officer: Julielynn Wong

Environmental control: Ambient with door open (door opened at 0930, re-opened at 1600, closed at 1900); no shade cover was used.

Average temperatures: 25 degrees Celsius

Low: 15 degrees Celsius

High: 32 degrees Celsius

Hours of supplemental light: None

Daily water usage for crops: 4 gallons

Water in Blue Tank—128 gallons

Time(s) of watering for crops: 1300; 1900

Changes to crops: None

Narrative: At 1600, the door was noted to be blown shut due to high winds. GHO suggests 3D printing a better door latch plus salad strainer, salad spinner, and fly swatter at MDRS.

Operations Report – March 27th

Crew 207 Operations Report 27-MAR-2019

SOL: 11

Name of person filing report: Dean Jin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): 14.5

Solar – SOC 36% @ 1825 (before David turned on generator for the night)

Diesel Reading – 50%

Propane Reading – 60%

Ethanol Free Gasoline – 2 gallons

Water (auxiliary tank) – 0 gallons (Note: Not in Use)

Water (static tank) – 180 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: 0

Water in GreenHab – 128 gallons

Static to Loft Pump used – Yes

Water Meter: 142162.4 gallons @ 1815

Toilet tank emptied: No

Deimos rover used: No, not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No

Hours: 85.1

Beginning charge: 100%

Ending charge: 100%

Currently charging: No – Charging Complete – Plugged In

Opportunity rover used: No

Hours: 54.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: No – Charging Complete – Plugged In

Curiosity rover used: No

Hours: 84.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: No – Charging Complete – Plugged In

Notes on rovers: Not Used

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) N/A

Reason for use: Not Used

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: None

HabCar used and why, where? Not Used

CrewCar used and why, where? Not Used

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: EV suits normal, all batteries read > 12.5 volts

Summary of Hab operations: Per recent electrical anomalies, checked SOC of electrical system as follows:

0800 SOC 70%, generator turned off by David

0900 SOC 60%, generator off

1100 SOC 60%, generator turned on by David

1200 SOC 60%, generator turned off by David

1600 SOC 46%, generator off

1825 SOC 36%, generator turned on by David

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: None (Note: observatory is closed)

Summary of health and safety issues: None

Sol Summary – March 27th

Crew 207 Sol Summary Report 27-MAR-2019

Sol: 11

Summary Title: Self-Replicating Replicators

Author’s name: Julielynn Wong

Mission Status: Nominal

Sol Activity Summary:

Crew 207 completed their data collection for the bike-powered 3D printing study. Today we used our bike-charged battery to successfully 3D print a tongue depressor out of food-safe, biodegradable plastic and a replacement printer LCD knob out of PETG. Results are now being compiled for remote analysis and manuscript preparation.

Crew 207 continued 3D printing in the Hab today for our thoracentesis training study.

The following items were identified as useful to 3D print at MDRS:

Filament disposal box

COM continues to fulfill GHO duties. XO continues to fulfill ENG duties.

COM will revise activity schedule for Crew 208 based on timelines observed with Crew 207.

Look Ahead Plan: Continue 3D printing in Hab. Will rehearse additional drone controller flight testing for Crew 208. Plan to visit Hanksville tomorrow (Sol 0 for Crew 208) to gather materials in preparation for arrival of Crew 208.

Anomalies in work: None

Weather: Sunny with high winds

Crew Physical Status: Nominal

EVA: n/a

Reports to be filed: Operations, Greenhab

Support Requested: None